[ r1 – r ]/a + [r2 – r]/b = …... options are c/R c^2/ (r3)^2 c/r3 c^2/r3

[ r1 – r ]/a  +  [r2 – r]/b = …...
 
options are
  1. c/R
  2. c^2/ (r3)^2
  3. c/r3
  4. c^2/r3

Grade:12th pass

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
(r1-r)/a+(r2-r)/b=
\frac{\frac{\Delta }{s-a}-\frac{\Delta }{s}}{a}+\frac{\frac{\Delta }{s-b}-\frac{\Delta }{s}}{b}
on simplifying we get
{}{\frac{\Delta }{s(s-a)}+{\frac{\Delta }{s(s-b)}}
{}{\frac{\Delta c}{s(s-a)(s-b)}
{}{\frac{\Delta c(s-c)}{\Delta ^{2}}
{}{\frac{c(s-c)}{\Delta}
c/r3
Hence 3 is the answer
Thanks and Regards
Shaik Aasif
askIITians faculty